OpenStudy (anonymous):

e^0.3x=0.3 i know you have to use ln but i keep getting stuck sorry i keep asking question but i have to pass a class and i cant miss any more. i have 10 questions left.

OpenStudy (jackellyn):

ln (natural log) both sides ln(e^(0.3x)) = ln(0.3) (0.3x)ln(e) = ln(0.3) 0.3x = ln(0.3) x = ln(0.3)/0.3 = -4.013